### v1

# Sense and Sensitivity: Examining the Influence of Semantic Recall on Long Context Code Reasoning

Abstract
Although modern Large Language Models (LLMs) support extremely large contexts, their effectiveness in utilizing long context for code reasoning remains unclear. This paper investigates LLM reasoning ability over code snippets within large repositories and how it relates to their recall ability. Specifically, we differentiate between lexical code recall (verbatim retrieval) and semantic code recall (remembering what the code does). To measure semantic recall, we propose SemTrace, a code reasoning technique where the impact of specific statements on output is attributable and unpredictable. We also present a method to quantify semantic recall sensitivity in existing benchmarks. Our evaluation of state-of-the-art LLMs reveals a significant drop in code reasoning accuracy as a code snippet approaches the middle of the input context, particularly with techniques requiring high semantic recall like SemTrace. Moreover, we find that lexical recall varies by granularity, with models excelling at function retrieval but struggling with line-by-line recall. Notably, a disconnect exists between lexical and semantic recall, suggesting different underlying mechanisms. Finally, our findings indicate that current code reasoning benchmarks may exhibit low semantic recall sensitivity, potentially underestimating LLM challenges in leveraging in-context information.
1 Introduction
Large Language Models (LLMs) are increasingly applied to industry coding tasks [15] that demand understanding of large codebases [25]. Recent advances [12, 46, 50] enable these models to process extremely long inputs, up to millions of tokens [1, 26, 44]. However, even state-of-the-art (SoTA) LLMs often fail to effectively utilize such long contexts for natural language reasoning [42]. They exhibit a ‘lost-in-the-middle’ (LIM) effect, where performance significantly degrades when relevant information is situated centrally within the input [39].
Though numerous benchmarks evaluate long-context capabilities for code [2, 3, 4, 52, 58], we are the first to investigate how the position of relevant code snippets within large irrelevant code context impacts LLM reasoning performance. Moreover, to understand the potential sources of this positional variance, we explore the relationship between models’ code reasoning abilities and their capacity to recall information from the input context. We introduce a key distinction between lexical code recall (retrieving code verbatim) and semantic code recall (remembering what the code does). While lexical recall can be measured directly through retrieval tasks [43, 18], measuring semantic recall is challenging with existing datasets, since failures in solving the downstream reasoning tasks are difficult to attribute to specific recall failures. Therefore, we propose SemTrace, a technique that remedies two major issues with attributing semantic recall in code reasoning. First, statements in SemTrace are randomized and thus hard to predict, which means the models cannot rely on pretrained knowledge or heuristics to solve the task instead of the input context. Second, statements in SemTrace independently affect the task outcome, removing the effect of error propagating through dependent statements. Lastly, to quantify semantic recall sensitivity in existing benchmarks, we propose a method based on measuring performance degradation on intentionally incomplete code.
Our comprehensive empirical evaluation across 10 SoTA LLMs, including GPT 4.1 and open-source models with up to 236B parameters, reveals several key findings. First, we observe a pronounced lost-in-the-middle effect in code reasoning tasks. This degradation is particularly severe on SemTrace, where accuracy for even top models can plummet by up to 94.07% when the target code is positioned mid-context. Second, we find that lexical recall performance strongly depends on retrieval granularity – most models excel at retrieving entire functions but struggle with line-by-line retrieval, suggesting they may only be capable of retrieving some information relative to the context of preceding lines. Third, we demonstrate a surprising disconnect between lexical and semantic recall capabilities, with several models strong in one dimension often weak in the other, indicating these abilities may rely on different underlying mechanisms. Finally, our analysis shows that current code reasoning benchmarks have relatively low semantic recall sensitivity, potentially underestimating the challenges LLMs face when reasoning about specific code details provided solely within the context.
In summary, our main contributions are: 1) introducing the critical distinction between lexical and semantic code recall, along with a rigorous method to quantify semantic recall sensitivity of existing benchmarks; 2) proposing SemTrace, a novel technique specifically designed to isolate and measure semantic code recall capabilities; 3) providing the first comprehensive analysis of the lost-in-the-middle effect in code reasoning tasks, revealing significant positional degradation and disconnects between different recall capabilities; 4) demonstrating that current code reasoning benchmarks may have low semantic recall sensitivity, highlighting the need for more targeted evaluation metrics to accurately assess LLMs’ code understanding abilities in long contexts.
2 Code Reasoning Framework
We focus on code reasoning tasks, where the goal is to understand code from different parts of a larger code repository in order to solve a specific problem such as code infilling, bug finding, or execution prediction. Consider a task reasoning function $R$ which captures a model’s reasoning capabilities on a given task. This is difficult to measure directly so it is empirically approximated using performance metrics on benchmarks representative of the task. Formally, $R:\mathcal{C}\rightarrow\mathbb{R}$, where $\mathcal{C}$ denotes the space of code snippets, and the output is a real-valued proxy for the model’s reasoning ability. For two tasks $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, $R^{\mathcal{T}_{1}}(\mathcal{C})>R^{\mathcal{T}_{2}}(\mathcal{C})$ means the model reasons about $\mathcal{C}$ better in $\mathcal{T}_{1}$ than $\mathcal{T}_{2}$.
In practical programming scenarios, a task-relevant code snippet of interest, denoted $C_{1}$, typically appears within a much larger code repository or context, denoted $C_{2}$, and is often difficult to precisely separate from $C_{2}$ [7, 24]. To model this, we define a mixing function $M$ (Equation 1) that describes how two independent code snippets $C_{1}$ and $C_{2}$ of length $n_{1}$ and $n_{2}$ tokens, respectively, combine to form a complete coding context. $Pos^{C_{1}}$ is a position vector of length $n_{1}+n_{2}$, where $C_{1}$ tokens’ positions are 1, and the rest are 0. Its complement $\overline{Pos^{C_{1}}}$ denotes the position vector corresponding to the tokens of $C_{2}$. $\otimes$ represents element-wise multiplication.
$M(C_{1},C_{2},Pos^{C_{1}})=\overline{Pos^{C_{1}}}\otimes C_{2}+Pos^{C_{1}}\otimes C_{1}$ (1)
Given the position of $C_{1}$ within the input context, we define two properties over the position $Pos^{C_{1}}$, where $Pos^{C_{1}}_{i}$ denotes its $i$-th element (Equation 2). These properties quantify how intertwined $C_{1}$ is with the surrounding code $C_{2}$. The first property is granularity, meaning the relative size of $C_{1}$ with respect to the full context. As the size of $C_{2}$ increases, the granularity of $C_{1}$ increases, making it a smaller with respect to the larger context. The second property is concentration, meaning the degree of spread of $C_{1}$ along the entire context window. Concentration takes a value of 1 when $C_{1}$ is a contiguous block of code and a value in the interval $(0,1)$ when $C_{1}$ is fragmented or non-contiguous.
$\underbrace{\frac{n_{2}}{n_{1}+n_{2}}}_{\text{granularity}}\qquad\underbrace{\frac{n_{1}}{max(\{i:Pos^{C_{1}}_{i}=1\})-min(\{i:Pos^{C_{1}}_{i}=1\})}}_{\text{concentration}}$ (2)
To better understand a model’s reasoning capabilities over a task-relevant code snippet $C_{1}$ when given alongside a task-irrelevant code snippet $C_{2}$, we propose a framework that captures four key degrees of freedom in the interaction between $C_{1}$ and $C_{2}$. Together, they form the complete input $S$ provided to the model. The model’s reasoning ability for a task $\mathcal{T}$, which requires understanding $C_{1}\in\mathcal{C}$ within the context $S=M(C_{1},C_{2},Pos^{C_{1}})$, is captured by the function $R^{\mathcal{T}}$ (Equation 3).
[…]
The terms in Equation 3 represent four distinct degrees of freedom. Terms (i) and (ii) correspond to the model’s understanding of individual code snippets when they are presented without interference from any other snippets. Term (iii) captures the influence ($Inf$) of $C_{2}$ on $C_{1}$, that is, how the presence of $C_{2}$ affects reasoning over $C_{1}$. This influence is expected to be positive for tasks where $C_{2}$ compliments $C_{1}$, such as few-shot learning examples [8], and negative for tasks where $C_{2}$ acts as a distraction or is orthogonal to $C_{1}$ [21, 42]. Finally, Term (iv) captures the impact of position – where $C_{1}$ appears in the input context – on reasoning performance. The average effect of position over $R^{\mathcal{T}}$ is represented by the expectation defined in Equation 4. Specifically, we marginalize the model’s reasoning performance over all possible code snippets at a given position $Pos^{C_{1}}$, where $Pr(C_{i},C_{j})$ denotes the joint probability of snippets $C_{i}$ and $C_{j}$ co-occurring in $\mathcal{C}$. In practice, this distribution is typically assumed to be uniform when evaluating over benchmark datasets. In this study, we leverage the variance in $R^{\mathcal{T}}$ due to positional effects (Equation 4) to investigate the model’s reasoning abilities. The precise relationship between these factors, represented by $F$, is unknown.
2.1 Prior Work on Code Reasoning
For a reasoning task, the terms $(i)$ and $(ii)$ in Equation 3 represent the task difficulty. When a task is too challenging for a model to solve in isolation, overall reasoning performance $R(C_{1}|S)$ is limited by the task difficulty $R(C_{1}\mid C_{1})$. In the context of code generation, this encapsulates the research on increasing task difficulty, such as docstring-based code generation [37] and input-output prediction [19]. Term $(iii)$ captures prior work on in-context learning in code reasoning [5, 9, 17, 32, 33] where the effect of $C_{2}$ compliments $C_{1}$. Conversely, this term also captures needle-in-a-haystack setups [38] where $C_{2}$ is designed as a distraction. Finally, the term $(iv)$ accounts for positional effects within the context. This includes the ‘lost-in-the-middle’ phenomenon in natural language reasoning [39] or ‘lost-at-the-end’ for vulnerability localization [49], demonstrating how the position of a relevant code snippet within a long context can significantly affect model performance.
2.2 In-Context Code Recall
Models tend to perform well on tasks that are well-represented in their training data. For instance, a model might be able to generate code for a well-known algorithm like quicksort more accurately than for algorithmically simpler but arbitrary code snippets. Effective reasoning over arbitrary or unfamiliar code snippets – especially those not seen during training – requires the ability to recall relevant parts of the code. In fact, the inability to recall specific parts of the code may render certain coding tasks impossible, as shown by the work on incremental consistency in code understanding [9].
However, prior work on in-context recall for LLMs often focuses on what we term Lexical Code Recall ($R^{L}$), defined as the ability to retrieve or reproduce code snippets from the input context verbatim, without necessarily understanding their meaning or behavior. This concept captures how well the model’s latent representation segments and preserves the surface form of the input: a model excelling solely at $R^{L}$ might operate by storing the one-hot representations of the input string segments. One can directly assess this ability through retrieval or cloze-style tasks. Modern LLMs excel at this [48], achieving near-perfect scores on the benchmarks testing these capabilities.
Orthogonal to lexical recall, we introduce Semantic Code Recall ($R^{S}$), defined as the ability to remember what a code snippet from the input context does. $R^{S}$ thus measures the usability of the model’s latent representation for performing downstream reasoning tasks that depend on the provided code. For instance, given an in-context implementation of quicksort and an unsorted array, $R^{L}(quicksort)$ corresponds to regenerating the quicksort code, whereas $R^{S}(quicksort)$ corresponds to leveraging the understanding derived from that code (the sorting logic) to correctly predict a sorted output array by tracing or simulating quicksort’s execution steps.
We further define the semantic recall sensitivity of a task. This property quantifies the degree to which successful task performance depends on $R^{S}$, as opposed to leveraging knowledge acquired during training. Consider an output prediction task. If the in-context function implements a standard algorithm like quicksort, a model might predict the sorted output using its pre-trained knowledge of sorting, even with imperfect semantic recall of the specific implementation details. Such a task exhibits low semantic recall sensitivity. Conversely, if the function implements a novel or arbitrary computation (e.g., returning a specific, non-standard permutation of the input array defined only within the context), the model must rely heavily on accurately recalling and interpreting the specific lines of code provided. This task would have high semantic recall sensitivity, as pre-trained knowledge alone is insufficient.
2.3 Measuring Semantic Code Recall
def f(x):
arr = [0, 0, 0, 0]
arr[0] = x - 43
arr[2] = x - 65
arr[1] = x + 88
arr[3] = x - 74
return arr
Listing 1: Example from SemTrace. Given input $x=81$, the model must predict the output list $[38,169,16,7]$ by correctly recalling and applying each assignment line.
Identifying semantic recall failures presents an underexplored challenge in prior work, as existing datasets suffer from two main issues: (a) the extent to which the model needs to rely on the input context to solve the task versus pretrained knowledge or heuristics is unclear and (b) most tasks involve a series of multiple mutually-dependent statements that determines the task outcome, rendering it difficult to attribute a task-level error to a failure to semantically recall a specific statement. To address this, we propose SemTrace, an output prediction task with high semantic recall sensitivity, where task-level errors can be directly attributed to failures in semantic recall of specific statements, rather than to general reasoning failures or deficiencies in pre-trained knowledge.
In SemTrace, the model receives a Python function that takes an integer $x$ as input. The function initializes a list and populates it through a series of assignment statements before returning the list. Crucially, each assignment modifies a distinct list element by adding a value $y$ (drawn uniformly from $[-100,99]$) to the input $x$. The order of these assignment lines within the function body is randomized. An example is shown in 1. Because the specific operations ($x+y$) follow no predictable pattern, the model can only rely on pretrained heuristics if it memorizes the solution to $x+y$ for all possible combinations of $x$ and $y$. Furthermore, since the probability of guessing a single operation correctly is $\frac{1}{200}$, for functions with $k$ assignments ($k$ drawn uniformly from $[4,10]$), the probability of guessing the entire output array by chance, even assuming knowledge of the generation process, is negligible (at most $(\frac{1}{200})^{4}=6.25\times 10^{-10}$). Correctly predicting the final array therefore necessitates accurate $R^{S}$ of all assignment lines, rendering SemTrace highly sensitive to $R^{S}$. Furthermore, since SemTrace’s lines are mutually independent and comparable in reasoning difficulty, we can measure the degree of $R^{S}$ by checking the number of array positions the LLM predicted correctly.
2.4 Measuring Semantic Recall Sensitivity
While SemTrace is designed a priori to exhibit high semantic recall sensitivity, it is also crucial to quantify this property for existing code reasoning tasks. Doing so allows us to understand the extent to which a task $\mathcal{T}$ requires a model to rely on semantic recall of the provided code snippet $C$, rather than leveraging potentially shortcut solutions based on pre-trained knowledge.
We propose estimating the semantic recall sensitivity by measuring the impact of withholding information from the code snippet. Specifically, we compare the model’s performance on the original task formulation with its performance in a scenario where the snippet is kept intentionally incomplete. Let $Inc(C)$ denote a set of incomplete versions $C^{inc}$ from the original snippet $C$ (e.g., by removing specific lines). The sensitivity of task $\mathcal{T}$ with respect to $C$, denoted $Sens^{\mathcal{T}}(C)$, can be estimated as the average normalized performance degradation across these incomplete versions (Equation 5).
$Sens^{\mathcal{T}}(C)=\frac{1}{|Inc(C)|}\sum_{C^{inc}\in Inc(C)}\frac{R^{\mathcal{T}}(C)-R^{\mathcal{T}}(C^{inc})}{R^{\mathcal{T}}(C)+\epsilon}$ (5)
Here, $R^{\mathcal{T}}(C)$ is the model’s reasoning performance on the original snippet, and $R^{\mathcal{T}}(C^{inc})$ is its performance on an incomplete version. The term inside summation captures the relative performance caused by the incomplete information in $C^{inc}\in Inc(C)$ (with $\epsilon>0$ added for numerical stability). A higher $Sens^{\mathcal{T}}(C)$ indicates that successfully performing task $\mathcal{T}$ on snippet $C$ heavily depends on semantically recalling the specific details within $C$, thus signifying high semantic recall sensitivity.
3 Empirical Relationship between Semantic Recall and Code Reasoning
This section investigates the relationship between LLMs’ code reasoning capabilities and their lexical and semantic recall abilities. We design controlled experiments manipulating the position of relevant code snippets and the overall input context size to understand how these two factors impact the performance of LLMs on code reasoning tasks.
Isolating Positional Effects. While realistic programming tasks often involve reasoning over code snippets that are highly dispersed throughout a large input context (i.e., $C_{1}$ and $C_{2}$ in Equation 3), we argue that such scenarios are inherently more difficult than reasoning over code where $C_{1}$ appears as a single, contiguous block clearly separated from irrelevant code $C_{2}$. In our setup, we consider the idealized case where $C_{1}$ is a contiguous code block with concentration equal to 1 (Equation 2). To further minimize the influence term $Inf(C_{2}\text{ on }C_{1})$, we ensure that $C_{2}$ is both unrelated to the task $\mathcal{T}$ and non-interfering with $C_{1}$. This setup resembles needle-in-a-haystack setups [38] where the irrelevant code $C_{2}$ functions as a "distractor". Under this controlled setting – with minimal influence and maximum concentration – we vary: (i) the position $Pos^{C_{1}}$ of the relevant code snippet across the input context to empirically estimate $\mathbb{E}(R^{\mathcal{T}},Pos^{C_{1}})$ as defined in Equation 4; and (ii) the granularity (Equation 2) of $C_{1}$ by increasing the length of the irrelevant snippet $n_{2}$, thereby expanding the coding context. This experimental design allows us to evaluate the model’s reasoning ability on task $\mathcal{T}$ under near-optimal conditions, isolating the effects of position and context length.
Models & Inference Setup. Our evaluation focuses on state-of-the-art LLMs, including top performers on coding benchmarks like EvalPlus [36], Copilot Arena [10], and SWE-Bench Verified [11, 25]. For some models, we also use multiple different sizes to analyze the impact of model scale within families. Specifically, we include GPT 4.1 (2025/04/14) [44], Qwen 2.5 Coder Instruct (7B and 32B) [22, 55], Llama 3 Instruct 3.1 8B and 3.3 70B [14], Gemma 3 It (12B and 27B) [16], DeepSeek Coder V2 Instruct (Lite and full-size) [20], and Codestral 22B v0.1 [51] (see § F for licensing information). Open-source models were served on a p5e.48xlarge AWS EC2 instance using vLLM [28] (see § I for details). In line with previous work [39] and to ensure reproducibility, we use greedy decoding for inference. To keep the GPT 4.1 costs manageable, we perform each experiment with GPT 4.1 only on one-eighth of each dataset, randomly sampled.
Prompting. We use query-aware contextualization [39] – placing the query both before and after the code – to enable decoder-only models to attend to the query when processing the code. Prompt templates are provided in § J, and an ablation study on query-aware contextualization is in § C.
3.1 How well do LLMs reason about code as the input context size increases?
We first evaluate how LLMs’ fundamental code understanding, crucial for downstream software engineering tasks (e.g., debugging, test generation), degrades as input context size increases. We measure accuracy on function-level input and output prediction using the CruxEval benchmark [19], which comprises 800 simple Python functions. Output prediction (CruxEval-O) probes execution tracing, while input prediction (CruxEval-I) assesses the ability to discern implicit assumptions.
Experimental Setup. Each CruxEval function is embedded within a context containing 20, 40, 60, or 80 distractor functions. The function’s position is systematically varied across 11 equally spaced locations within this context. Distractor functions are randomly selected Python functions drawn from CodeSearchNet [40, 23], a large and popular dataset. These functions are filtered by character count (25th–75th percentile) to exclude length outliers.
Results and Discussion. Results (Figure 1) reveal a pronounced LIM effect for both input and output prediction. Accuracy decreases significantly as the target function moves towards the context’s fourth quintile and as the number of distractors increases. Relative losses are often steeper for smaller models (e.g., Qwen 2.5 Coder 7B’s input accuracy decreases by up to 79.55%, DeepSeek Coder V2 Lite’s output accuracy by up to 74.58% from their respective maximums). Conversely, while larger models achieve higher peak accuracy, they often suffer larger absolute performance drops. For example, GPT 4.1 and DeepSeek Coder V2 drop up to 15.00 and 30.00 percentage points (pp) on input and 13.00 and 33.75 pp on output prediction, respectively. See § E for more details.
The Llama 3 family, particularly Llama 3.1 8B, shows greater positional robustness, though it still experiences significant relative drops (up to 46.67% for input prediction, 31.63% for output prediction). Crucially, this degradation occurs well within models’ advertised context limits (GPT 4.1 supports up to 1M input tokens), demonstrating that theoretical support for large context windows does not guarantee effective utilization, especially for code located far from context boundaries.
Figure 1: Input and output prediction accuracy on CruxEval, relative to the maximum accuracy achieved by each model for a given number of distractor functions, by position of the target function within the input context and the number of distractors. Average distractor is $\sim 200$ tokens.
3.2 What is the effect of increasing input context size on LLMs’ lexical code recall ability?
To explain the code reasoning patterns observed in § 3.1, we first investigate the models’ lexical code recall ($R^{L}$) ability. We prompt models to retrieve the target function ($C_{1}$) from the context ($C_{2}$), testing both function-level (entire function body) and line-level (single line) granularity.
Experimental Setup. We use identical input contexts from § 3.1. To facilitate precise line identification, we prepend each line with a unique 6-digit hexadecimal key, framing the task similar to key-value retrieval [30, 39, 45, 58]. We found that this method yields better results than line numbering or no reference, consistent with [30], without affecting the models’ code reasoning performance (see § D). Success on this task requires an exact match with the original line or function body.
Results and Discussion. Our findings reveal a clear dichotomy in performance (Figure 2). Most models perform well on function-level retrieval across positions and context sizes, though positional sensitivity is noticeable in smaller models (Llama 3.1 8B, DeepSeek Coder V2 Lite, Qwen 2.5 Coder 7B). However, line-level retrieval exhibits a strong LIM pattern as the context size increases and the target function approaches the context’s fourth quintile, with a 16.05% average accuracy decrease across models at 20 distractors and up to 36.42% drop at 80 distractors. Only GPT 4.1, Gemma 3 27B, and Qwen 2.5 Coder 32B maintain stable line-level performance. This disparity suggests that while LLMs are capable of retrieving larger, coherent code blocks such as entire functions, many struggle with fine-grained, line-level retrieval. This may be because some lines are easily predictable from the context of preceding lines. Therefore, certain models might only recall these lines with respect to previous lines – in an autoregressive manner – especially in long contexts. Notably, the Qwen 2.5 Coder family shows relative consistency in these lexical recall tasks across model sizes, implying factors beyond just scale influence positional robustness for lexical recall.
Figure 2: Lexical recall accuracy at function- and line-level granularity on CruxEval against different target function positions within the input context for varying sets of distractor functions.
3.3 How does increasing input context size affect LLMs’ semantic code recall ability?
We evaluate LLMs’ semantic recall ability ($R^{S}$) using SemTrace (defined in § 2.3), reporting accuracy over exact and partial matches with the expected output. The former metric captures the performance on a task highly sensitive to semantic recall, whereas the latter is a more fine-grained measure of $R^{S}$.
Experimental Setup. We generate 800 SemTrace examples and evaluate output prediction using the same distractor setup as in § 3.1. For performance without distractors, see § A. Partial match accuracy is computed by parsing the model’s output and computing the share of correct elements in the list.
Results and Discussion. Figure 3 reveals dramatic differences on SemTrace. Despite SemTrace’s simplicity, smaller models such as Gemma 3 12B and Llama 3.1 8B achieve near-zero accuracy even with only 20 distractors, rendering the task effectively impossible for them. In contrast, larger models (GPT 4.1, DeepSeek Coder V2, Llama 3.3 70B, Qwen 2.5 Coder 32B) achieve 36.38% higher average accuracy on SemTrace than on CruxEval-O with 20 distractors. However, the accuracy of these models except for GPT 4.1 plummets on average by up to 83.23% or 54.38 pp as the context size increases and the target function moves towards the middle/latter half of the context. We also find that several models including Llama 3.3 70B generate answers like $81-43$ instead of $38$ more frequently as the target function approaches the context’s center (explored in § H). This extreme positional sensitivity, isolated by SemTrace, underscores the fragility of semantic recall. While performance is often robust at the context start/end (DeepSeek Coder V2 being an exception, degrading at the start of longer contexts), the overall U-shaped pattern mirrors that seen in CruxEval-O (Figure 1), suggesting shared underlying limitations related to semantic recall over long contexts.
A notable exception is GPT 4.1, which performs nearly perfectly on SemTrace. We hypothesize that GPT 4.1 has memorized 2-digit addition/subtraction, thus performing retrieval rather than reasoning – a task where we have shown no LIM effect occurs. To test this hypothesis, we additionally ran GPT 4.1 on SemTrace with 3-, 4-, 5-, and 6-digit addition/subtraction. As shown in Figure 4, GPT 4.1 performs consistently well on 2- and 3-digit SemTrace, but the LIM effect reemerges for higher digits, with accuracy dropping up to 31.00 pp on 4-digit SemTrace and up to 42.65% on 6-digit SemTrace. The LIM effect also intensifies as the number of SemTrace digits increases.
Intriguingly, we find little correlation between lexical (§ 3.2) and semantic recall across models. Models strong lexically (e.g., Gemma 3 27B, Qwen 2.5 Coder 32B) degrade significantly here, while Llama 3.3 70B shows relative stability despite line-level lexical recall weaknesses. This suggests potentially distinct mechanisms govern lexical and semantic recall that vary across models and scales.
Figure 3: Semantic recall accuracy using exact and partial matches on SemTrace against different target example positions within the input context for varying sets of distractor functions.
Figure 4: GPT 4.1’s acccuracy on $k$-digit SemTrace with 80 distractors, relative to its maximum accuracy achieved for each digit.
3.4 How sensitive is CruxEval to semantic code recall?
The smoother LIM effect observed for CruxEval (§ E) contrasting with the sharp drop on SemTrace (§ 3.3) when comparing absolute accuracies suggests CruxEval might be less demanding on semantic recall as defined in § 2.2. To quantify this, we perform a counterfactual experiment removing lines from CruxEval functions to measure how information loss impacts performance (method described in § 2.4).
Experimental Setup. For each CruxEval function, we generate variants by removing all possible combinations of lines (preserving the function signature), creating 71,994 incomplete functions. We evaluate LLM input/output prediction accuracy on these functions without distractors, aggregating results by the percentage of lines removed.
Results and Discussion. If models truly reason about the input context – leveraging semantic recall – we would expect to see a sharp, exponential decay in accuracy as more lines are removed, similar to Python interpreter’s performance (detailed in § B). Conversely, models relying primarily on pretrained knowledge and heuristics should show more gradual accuracy decay. Using non-linear least squares, we fit exponential trend lines to the aggregated results in Figure 5. Most tested LLMs demonstrate remarkable resilience to missing information on CruxEval, particularly for input prediction. Removing 50% of lines causes only a 44.15% accuracy drop for input prediction and 59.74% for output prediction. For output prediction, larger models (GPT 4.1, DeepSeek Coder V2, Llama 3.3 70B, Qwen 2.5 Coder 32B) show sharper performance declines than smaller models, suggesting they rely more on semantic recall than on learned heuristics and pattern-matching. This result further explains their superior performance on SemTrace compared to smaller models. Moreover, these findings indicate that tasks like SemTrace, which are highly sensitive to semantic recall, may better assess true in-context reasoning abilities while minimizing pretrained heuristics.
Figure 5: Sensitivity of CruxEval input/output prediction to line removal. We fit exponential trend lines using non-linear least squares and leverage bootstrapping to compute 95% confidence intervals.
4 Related Work
As language models grow in size and capability, extending their effective context length has become a major research focus. Prior work on long-context language models either focuses on modifying positional encoding [46, 47, 50] or reworking the attention mechanism [6, 13, 35, 57]. To understand where these methods fall short, others have proposed specialized benchmarks [21, 34, 38] and stress tests [29, 39] to evaluate the limitations of long-context models. More recently, attention has shifted to long-context reasoning through novel benchmarks [27, 56, 58] and specialized training and inference strategies aimed at bolstering reasoning performance over long context [31, 53, 54]. A growing body of research examines different factors in Equation 3. We cover these works in § 2.1.
5 Limitations and Future Work
In our evaluation of positional effects, we only varied granularity but did not explore concentration of relevant code snippets (as defined in Equation 2). Further investigation is required to pinpoint the underlying factors contributing to performance degradation associated with positional variance over longer contexts. Additionally, we did not evaluate reasoning models due to their significantly different inference architectures, which complicate isolating positional effects. These models are also computationally expensive to evaluate, placing them beyond the scope of this work.
6 Conclusion
In this paper, we investigated the reasoning capabilities of LLMs over code snippets within large repositories, providing a distinction between lexical code recall and semantic code recall. Our evaluation shows that LLMs’ accuracy at performing code reasoning drops when the relevant information is located centrally within the context. This performance degradation is particularly exacerbated by our technique, SemTrace, which requires an LLM to exhibit high semantic recall. In addition, we also uncover that LLMs struggle with finer-granularity, line-level retrieval as compared to function-level retrieval, and a clear disconnect between their different recall capabilities. Finally, we show that existing benchmarks may have low semantic recall sensitivity, underscoring the need for benchmarks that can rigorously test semantic recall capabilities.
Broader Impact Statement
This paper highlights how the positional variance of code within context affects reasoning tasks, advancing our understanding of LLMs’ reasoning capabilities in long-context scenarios. While there may be broader societal implications, we do not consider them significant enough to emphasize here.
Appendix A Baseline Performance (without Distractors)
Table 1: Baseline performance (accuracy) of LLMs on CruxEval Input Prediction (CruxEval-I), CruxEval Output Prediction (CruxEval-O), and SemTrace Output Prediction (SemTrace).
Model CruxEval-I CruxEval-O SemTrace
Codestral 22B 44.62 50.38 23.25
GPT 4.1 (2025/04/14) 79.00 77.00 100.00
Deepseek Coder V2 Lite 20.00 21.00 4.12
Deepseek Coder V2 36.50 54.00 96.00
Gemma 3 12B 39.50 19.25 15.88
Gemma 3 27B 50.62 25.87 12.25
Llama 3.1 8B 19.50 32.62 16.00
Llama 3.3 70B 18.25 56.75 87.12
Qwen2.5 Coder 7B 31.62 46.62 22.25
Qwen2.5 Coder 32B 45.50 61.75 85.25
Baseline performance results for all tested open-source models are presented in Table 1. Most models achieved higher accuracy on CruxEval-O than CruxEval-I. The Gemma 3 family was a notable exception, demonstrating outstanding accuracy on CruxEval-I but low accuracy on CruxEval-O. On SemTrace, the largest models tested (GPT 4.1, DeepSeek Coder V2, Llama 3.3 70B, Qwen 2.5 Coder 32B) attained remarkably high accuracy, while smaller models showed dramatically lower accuracy.
Appendix B Python Interpreter’s Sensitivity to Line Removal
To understand the strict line redundancy in CruxEval output prediction (CruxEval-O), we perform output prediction by running the Python interpreter on all of the incomplete functions from § 3.4. As shown in Figure 6, we find a sharp exponential decay in accuracy, where removing only one line leads to a sharp, 76.14 percentage point drop in accuracy, and the output prediction accuracy approaches zero as more than 20% of lines is removed.
Figure 6: Sensitivity of CruxEval output prediction (CruxEval-O) to line removal, as performed by the Python interpreter. We fit exponential trend lines using non-linear least squares and leverage bootstrapping to compute 95% confidence intervals.
Appendix C Ablation Study: Query-Aware Contextualization
Figure 7: Input and output prediction accuracy on CruxEval without query-aware contextualization for a given number of distractor functions by position of the target function within the input context and the number of distractors.
Figure 8: Lexical recall accuracy at function- and line-level granularity on CruxEval without query-aware contextualization against different target function positions within the input context for varying sets of distractor functions.
While, following [39], our main experiments employ query-aware contextualization (QAC), we evaluate the impact of its removal here. We re-ran input/output prediction and function-/line-level retrieval tasks without QAC on a 100-example subset of CruxEval for all open-source models we tested.
The effect of removing QAC on input/output prediction is inconsistent (Figure 7). Models tend towards slightly improved input prediction accuracy without QAC, but decreased output prediction accuracy. The Lost-in-the-Middle (LIM) effect’s magnitude also changes unpredictably: it increases for some models without QAC, while decreasing for others. Performance variations across target positions also appear greater without QAC.
Conversely, removing QAC substantially hinders retrieval performance (Figure 8). With QAC, most models approach perfect function-level retrieval; without it, they exhibit a noticeable LIM effect. For line-level retrieval, models like Qwen 2.5 Coder 32B and Gemma 3 27B eliminate the LIM effect with QAC, but without QAC, all models suffer from a more severe LIM effect than they do with QAC present. These results indicate that QAC aids lexical recall and provides a significant retrieval advantage, particularly for longer input contexts. This ablation also serves as additional evidence of LLMs using distinct mechanisms for code reasoning and lexical recall.
Appendix D Ablation Study: Effect of Prepending Lines with Hexadecimal Keys on Code Understanding
To understand the effect of prepending lines with hexadecimal keys on the model’s ability to reason about code, we measure the open-source models’ performance on CruxEval. Based on the results (Table 2), we find that the tested models’ ability to perform these tasks is not significantly affected.
Table 2: Performance (accuracy) of LLMs on CruxEval Input Prediction (CruxEval-I) and CruxEval Output Prediction (CruxEval-O) with 6-digit hexadecimal keys prepended to each line.
Model CruxEval-I CruxEval-O
Codestral 22B 43.75 50.38
Deepseek Coder V2 Lite 20.88 18.88
Deepseek Coder V2 28.38 54.75
Gemma 3 12B 36.88 19.38
Gemma 3 27B 46.63 23.50
Llama 3.1 8B 18.38 30.38
Llama 3.3 70B 20.00 57.25
Qwen2.5 Coder 7B 29.38 42.13
Qwen2.5 Coder 32B 44.50 61.38
Appendix E Performance on CruxEval in Terms of Absolute Accuracy
In Figure 9, we additionally provide the results of CruxEval input prediction (CruxEval-I) and output prediction (CruxEval-O) in terms of absolute accuracy.
Figure 9: Input and output prediction accuracy on CruxEval for a given number of distractor functions by position of the target function within the input context and the number of distractors.
Appendix F Model and Dataset Licenses
We include the licenses for models and datasets in Table 3.
Table 3: Licensing information for the evaluated models and employed datasets.
Model Name License Codestral 22B MNLP-0.1 GPT 4.1 Proprietary DeepSeek Coder V2 Deepseek-license DeepSeek Coder V2 Lite Deepseek-license Llama 3.1 8B Llama3.3 Llama 3.3 70B Llama3.3 Qwen 2.5 Coder 32B Apache-2.0 Qwen 2.5 Coder 7B Apache-2.0 Gemma 3 12B Gemma Gemma 3 27B Gemma Dataset Name License CruxEval MIT License CodeSearchNet MIT License
Appendix G Architectural Choices of Models
Table 4 includes a detailed overview of the architectural choices across the evaluated language models, highlighting differences in positional encoding, attention mechanisms, and context length.
In general, larger models tend to outperform their smaller counterparts within the same family in all tasks, with the exception of the function retrieval task (Figure 2), where most models demonstrate strong performance regardless of size. Aside from GPT 4.1, Qwen 2.5 Coder 32B demonstrates superior performance across most tasks, despite having fewer parameters than DeepSeek Coder V2 236B and Llama 3.3 70B. In contrast, DeepSeek Coder V2 Lite consistently ranks among the lowest-performing models evaluated, while having 15.7B parameters which is more than Llama 3.1 8B, Qwen 2.5 Coder 7B, and Gemma 3 12B. These results confirm that while increasing the model size improves the performance in the same family, it does not guarantee better long context performance across different architectures.
DeepSeek Coder V2 and DeepSeek Coder V2 Lite are the only models that utilize mixture of experts while not using grouped attention query. On both SemTrace (Figure 3) and input/output prediction (Figure 1), DeepSeek Coder V2 shows more extreme positional swings than the other models—its accuracy plunges further in the middle and then surges back at the end. However, it shows relatively lower variance in retrieval tasks compared to its smaller counterpart. On the other hand, DeepSeek Coder V2 Lite shows more variance in retrieval tasks (Figure 2), especially as the number of distractors increases, while performing worse and with less variance in the other tasks. Additionally, with 60–80 distractors, DeepSeek Coder V2 exhibits an end-position advantage in output prediction and a start-position advantage in input prediction; the former is also evident in SemTrace.
Gemma 3 27B performs nearly perfectly in lexical recall; however, in SemTrace it performs worse than larger models (DeepSeek Coder V2, Qwen 2.5 Coder 32B, Llama 3.3 70B). While this behavior could be due to sliding window or positional encoding types which are both different from all other models, it is not possible to draw definitive conclusions based on the current data. Other variables – such as the training procedure, the composition of the dataset, etc. – could also have played a significant role in the model’s superior performance in lexical recall. This observation underscores that an LLM’s semantic recall ability does not necessarily align with its lexical recall performance.
Men et al. [41] introduced a theory that establishes the absolute minimum base frequency value required to achieve a given context-length capability in RoPE. For our tasks, the maximum context spans approximately 31k tokens, which corresponds to a minimum base RoPE frequency of roughly 861k. Codestral 22B, Qwen 2.5 Coder 32B, and Qwen 2.5 Coder 7B models exceed this requirement by using a base frequency of 1M. Moreover, it is noteworthy that none of our tasks surpass the documented context length of Codestral 22B– the only model among these with a 32k token window – while the remaining models provide a 128k/1M token window.
Table 4: Architectural configurations of various language models evaluated in this study, including their positional encoding types. MoE: Mixture of Experts, GQA: Grouped Query Attention. RoPE: Rotary Position Embedding, YaRN: Yet another RoPE extensioN.
Model Name Positional Encoding MoE Sliding Window GQA Context Length Codestral 22B RoPE (Theta: 1M) No No Yes 32768 DeepSeek Coder V2 YaRN (Theta: 10k) Yes No No $\sim$128k DeepSeek Coder V2 Lite YaRN (Theta: 10k) Yes No No $\sim$128k Llama 3.1 8B Llama3 RoPE (Theta: 500k) No No Yes 128k Llama 3.3 70B Llama3 RoPE (Theta: 500k) No No Yes 128k Qwen 2.5 Coder 32B RoPE (Theta: 1M) No No Yes $\sim$128k Qwen 2.5 Coder 7B RoPE (Theta: 1M) No No Yes $\sim$128k Gemma 3 12B Linear RoPE (Scaled by 8.0, Theta: 1M) No 1024 Yes 128k Gemma 3 27B Linear RoPE (Scaled by 8.0, Theta: 1M) No 1024 Yes 128k GPT 4.1 - - - - 1M
Appendix H Resolution Errors in SemTrace
Despite explicitly prohibiting partial results in the prompt, we find that several models still at times generate answers such as $81-43$ instead of $38$. Since these expressions are interpreted by Python as equivalent, we don’t penalize such responses. However, our supplementary analysis (Figure 10) reveals that especially Gemma and Llama family models provide these "unresolved" outputs much more frequently when the target function appears near the middle of the input context. We hypothesize that this represents an attempt to circumvent semantic recall limitations through lexical recall.
Figure 10: Number of unresolved outputs on SemTrace against different target example positions within the input context for varying sets of distractor functions.
Appendix I Detailed Setup Information
Machine Details. We performed all experiments using AWS EC2 p5e.48xlarge instance equipped with 192 cores, 2048GB RAM, and eight NVIDIA H200 GPUs on Ubuntu 22.04 with CUDA 12.4. To serve open-source LLMs, we use vllm 0.7.1 [28]. We access GPT 4.1 through OpenAI API.
Execution Time. For the final evaluation runs, we spent total of 47 GPU-days on open-source model inference. We have spent about 2 days running experiments on GPT 4.1 through OpenAI API. We estimate that total usage, including reruns and development, might be 1.5-2 times higher than our evaluation runs.
Appendix J Prompt Templates
In this section we include templates for input prediction (Figure 11), output prediction (Figure 12), function-level retrieval (Figure 13), and line-level retrieval (Figure 14) tasks. We leverage assistant prefill such that each model provides a predictable and easy-to-parse response.
User:
You are given a number of Python functions and an assertion containing an output of one of the functions. Find any input such that executing that function on the input leads to the given output. There may be multiple answers, but you should only output one.
[ASSERTION]
assert {output} == f(??)
[/ASSERTION]
[FUNCTIONS]
{code_block}
[/FUNCTIONS]
[ASSERTION]
assert {output} == f(??)
[/ASSERTION]
You are given a number of Python functions and an assertion containing an output of one of the functions. Find any input such that executing that function on the input leads to the given output. There may be multiple answers, but you should only output one.
Assistant:
Sure! Here is the corresponding input:
‘‘‘python
assert {output} == f(
Figure 11: Input Prediction Chat Completion Prompt Template.
User:
You are given a number of Python functions and an assertion containing an input to one of the functions. Complete the assertion with a literal (no unsimplified expressions, no function calls) containing the output when executing the provided code on the given input, even if the function is incorrect or incomplete. Do NOT output any extra information.
[ASSERTION]
assert f({input}) == ??
[/ASSERTION]
[FUNCTIONS]
{code_block}
[/FUNCTIONS]
[ASSERTION]
assert f({input}) == ??
[/ASSERTION]
You are given a number of Python functions and an assertion containing an input to one of the functions. Complete the assertion with a literal (no unsimplified expressions, no function calls) containing the output when executing the provided code on the given input, even if the function is incorrect or incomplete. Do NOT output any extra information.
Assistant:
Sure! Here is the corresponding output:
‘‘‘python
assert f({input}) ==
Figure 12: Output Prediction Chat Completion Prompt Template.
User:
Each line in the code block below starts with a random key. I’m looking for a function starting at key ‘{start}‘ and ending at key ‘{end}‘ in the code snippet below. Can you help me find it?
‘‘‘python
{code_block}
Each line in the code block above starts with a random key. I’m looking for a function starting at key ‘{start}‘ and ending at key ‘{end}‘ in the code snippet above. Can you help me find it?
Assistant:
Sure! Here is the full function starting at key ‘{start}‘ and ending at key ‘{end}‘:
‘‘‘python
Figure 13: Function-level Retrieval Chat Completion Prompt Template.
User:
Each line in the code block below starts with a random key. I’m looking for a line with key ‘{key}‘ in the code snippet below. Can you help me find it?
‘‘‘python
{code_block}
Each line in the code block above starts with a random key. I’m looking for a line with key ‘{key}‘ in the code snippet above. Can you help me find it?
Assistant:
Sure! Here is the line with key {key}:
‘‘‘python
Figure 14: Line-level Retrieval Chat Completion Prompt Template.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims made in the abstract and introduction are accurately laid down and supported by the extensive empirical evaluation results presented in § 3.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of our current approach are specified in § 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: No theoretical results are presented in the paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Our experimental setup is detailed in § 3 with our open-sourced code repo containing additional instructions to reproduce our reported results from scratch.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our framework is open-sourced and has a detailed readme to reproduce the results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The test details for each experiment are detailed in § 3 and in the accompanying code repository.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Our results are deterministic since we only infer from the trained models using greedy decoding. Consequently, error bars do not apply to our lexical and semantic recall experiments. The only exception is our semantic recall sensitivity experiment (§ 3.4) where we report 95% bootstrapped confidence intervals.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We list the compute resources used in § 3 with further details about the setup and time taken to run the different experiments in § I.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: No human subjects were involved in this work. All datasets and models are cited along with their licenses attached in § F.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The broader impacts of our study are discussed in Broader Impact Statement.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No such artifacts are released.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All models and datasets are appropriately cited and their licensing information can be found in § F.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The experimental setup detailed in § 3, coupled with detailed instructions in our open-source repository, includes sufficient information to reproduce our reported results.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No research with human subjects or crowdsourcing experiments were performed as part of this work.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No research with human subjects or crowdsourcing experiments were performed as part of this work.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 Figure 14: Line-level Retrieval Chat Completion Prompt Template.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The claims made in the abstract and introduction are accurately laid down and supported by the extensive empirical evaluation results presented in § 3.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: The limitations of our current approach are specified in § 5.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: No theoretical results are presented in the paper.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: Our experimental setup is detailed in § 3 with our open-sourced code repo containing additional instructions to reproduce our reported results from scratch.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: Our framework is open-sourced and has a detailed readme to reproduce the results.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The test details for each experiment are detailed in § 3 and in the accompanying code repository.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34.
-Justification: Our results are deterministic since we only infer from the trained models using greedy decoding.
-Consequently, error bars do not apply to our lexical and semantic recall experiments.
-The only exception is our semantic recall sensitivity experiment (§ 3.4) where we report 95% bootstrapped confidence intervals.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We list the compute resources used in § 3 with further details about the setup and time taken to run the different experiments in § I.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: No human subjects were involved in this work. All datasets and models are cited along with their licenses attached in § F.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: The broader impacts of our study are discussed in Broader Impact Statement.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: No such artifacts are released.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: All models and datasets are appropriately cited and their licensing information can be found in § F.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: The experimental setup detailed in § 3, coupled with detailed instructions in our open-source repository, includes sufficient information to reproduce our reported results.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: No research with human subjects or crowdsourcing experiments were performed as part of this work.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: No research with human subjects or crowdsourcing experiments were performed as part of this work.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.